OpenStudy (anonymous):

Really need help with algebra, been stuck on this, please help . Jean and Mark are going to fill a pool with 2 different sized hoses. Jean can fill the pool in 8 hours, while Mark can complete it in 12 hours. Their supervisor thinks that the job will take 10 hours to complete if they work together. Explain each step in solving this equation and determine if the supervisor is correct or not.

OpenStudy (anonymous):

I think I know how to set up the equation

OpenStudy (anonymous):

\[\frac{ 1 }{ 8 } + \frac{ 1 }{ 12 } = \frac{ 1 }{ x } \] x would be how many hours they could do it together, and since we are trying to prove she's correct we don't really do anything with the 10 right now?

OpenStudy (anonymous):

Prove if she's correct*

OpenStudy (solomonzelman):

Jean = 8h Mark = 12h supervisor=10h Answer this, is it given that Jean=8h and Mark=8h is correct?

OpenStudy (anonymous):

Yes, it's given.

OpenStudy (solomonzelman):

(I meant mark=12h)

OpenStudy (anonymous):

and I knew what you meant no worries :)

OpenStudy (solomonzelman):

So the supervisor says that it would take them 10h. know if they need 10 hours each (in average to complete the job), the supervisor is incorrect because if they work together they will finish twice as fast, or in 5 hours (not 10).

OpenStudy (anonymous):

Wow thanks :)

OpenStudy (solomonzelman):

It's just a common sense question, no formula required.

OpenStudy (anonymous):

btw, do you understand this question? George tells you that when variables are in the denominator, the equation becomes unsolvable. "There is a value for x that makes the denominator zero, and you can't divide by zero," George explains. Using complete sentences, demonstrate to George how the equation is still solvable.

OpenStudy (solomonzelman):

I think so.

OpenStudy (anonymous):

Because I don't really get exactly what they're asking.

OpenStudy (solomonzelman):

Lets rephrase the question. Prove that the equation a has solution when x is in the denominator.

OpenStudy (anonymous):

So it's kinda the stuff we did last night? When a few of them had a denominator of x?